OpenStudy (anonymous):

A rectangular prism has a volume of 48 ft 3. If the height of the prism is 4 ft, and the length is 3 feet, what is the width? 6 ft 4 ft 3 ft 2 ft

Directrix (directrix):

The volume of a rectangular prism is l*w*h. V = l*w*h and 48 = 4*3*w So, solve 48 = 12 w to get the width